OpenStudy (anonymous):

Please Help ! Ill fan you :) What is the point-slope form of the equation of a line that passes through the point (6, -8 ) and has a slope of -2 ? A. y + 8 = 2(x - 6 ) B. y + 8 = -2(x - 6 ) C. y - 6 = -2(x - 8 ) D. y - 8 = -2(x + 6 )

OpenStudy (justin_lee):

Point slope form looks like this: \[y - y_{1} = m ( x - x_{1})\] You know your y1 is -8 and your x1 is 6. You know your slope is m = 2. Let's plug it in: Left hand side: \[y - y_{1} = y - (-8) = y + 8\] Right hand side: \[m(x - x_{1}) = -2(x - 6)\] Equating both sides gives: \[y+8=-2(x-6)\] Which should be the correct answer then?

OpenStudy (anonymous):

B ? @Justin_Lee

OpenStudy (justin_lee):

Yes.

OpenStudy (justin_lee):

As long as the info you give at the beginning has no typo.

OpenStudy (anonymous):

no theres no typo